what is the average of 3/4 and 5/6?
Found 2 solutions by rfer, ewatrrr:
Answer by rfer(16322)

(3/4+5/6)/2=
(9/12+10/12)/2=
(19/12)/(2/1)=
19/12*1/2=19/24

Answer by ewatrrr(24785)